# The Z-Gromov-Wasserstein Distance

### v1

@@ Abstract
The Gromov-Wasserstein (GW) distance is a powerful tool for comparing metric measure spaces which has found broad applications in data science and machine learning. Driven by the need to analyze datasets whose objects have increasingly complex structure (such as node and edge-attributed graphs), several variants of GW distance have been introduced in the recent literature. With a view toward establishing a general framework for the theory of GW-like distances, this paper considers a vast generalization of the notion of a metric measure space: for an arbitrary metric space $Z$, we define a $Z$-network to be a measure space endowed with a kernel valued in $Z$. We introduce a method for comparing $Z$-networks by defining a generalization of GW distance, which we refer to as $Z$-Gromov-Wasserstein ($Z$-GW) distance. This construction subsumes many previously known metrics and offers a unified approach to understanding their shared properties. The paper demonstrates that the $Z$-GW distance defines a metric on the space of $Z$-networks which retains desirable properties of $Z$, such as separability, completeness, and geodesicity. Many of these properties were unknown for existing variants of GW distance that fall under our framework. Our focus is on foundational theory, but our results also include computable lower bounds and approximations of the distance which will be useful for practical applications.
@@ 1 Introduction
where the infimum is over probability measures $\pi$ on $X\times Y$ whose left and right marginals are $\mu_{X}$ and $\mu_{Y}$, respectively. Intuitively, such a measure describes a probabilistic correspondence between the sets $X$ and $Y$, and the integral in (1) measures the extent to which such a correspondence distorts the metric structures of the spaces. The optimization problem (1) therefore seeks a probabilistic correspondence which minimizes total metric distortion. It was shown in [44] that the Gromov-Wasserstein $p$-distance defines a metric on the space of all triples $(X,d_{X},\mu_{X})$, considered up to measure-preserving isometries, when $(X,d_{X})$ is compact and $\mu_{X}$ is fully supported (these assumptions can be relaxed [49, 67]).
The Gromov-Wasserstein (GW) framework has become a popular tool in data science and machine learning—see [22, 21, 14, 76, 20, 1], among many others—due to its flexibility in handling diverse data types, its robustness to changes in metric or measure structures [53, Theorem 5.1], and recent advances in scalable and empirically accurate computational schemes for its estimation [56, 77, 18, 60, 40, 73]. It was observed in [56] that the formula for GW distance in (1) still makes sense when the condition that $d_{X}$ and $d_{Y}$ are metrics is relaxed; that is, the formula gives a meaningful comparison between structures of the form $(X,\omega_{X},\mu_{X})$, where $\omega_{X}:X\times X\to\mathbb{R}$ is an arbitrary measurable function. This point of view was later studied formally in [19], where it was shown that the GW distance defines a metric on the space of such triples $(X,\omega_{X},\mu_{X})$, considered up to a natural notion of equivalence. This is convenient, for example, when handling graph datasets, where it is natural to represent a graph’s structure through its adjacency function, Laplacian, or heat kernel.
Driven by applications to machine learning on increasingly complex data types, several variants of GW distance have been introduced in the literature [45, 72, 35, 48, 2, 78]. For example, [72] introduces an adaptation of GW distance which is equipped to handle graphs endowed with node features (i.e., whose nodes are endowed with values in some auxiliary metric space); this was recently extended in [78] to a framework treating graphs with edge features as well. Each time that a new variant of GW distance is introduced in the literature, its metric properties have been re-established. Overviewing these proofs reveals that they tend to follow a common template, suggesting the existence of a higher-level explanation of their shared properties.
@@ 1.1 Main Contributions and Outline
  • Theorem 29 says that, when $(Z,d_{Z})$ is separable, the $Z$-GW distance $\mathrm{GW}_{p}^{Z}$ defines a metric on the space $\mathcal{M}_{\sim}^{Z,p}$ consisting of $Z$-networks, considered up to a natural notion of equivalence (Definition 28). As a consequence, Corollary 30 shows that Fused GW and Fused Network GW distances are metrics; these were previously only shown to satisfy a certain weak triangle inequality. The proof of Theorem 29 relies on a technical result, which says that the solution of the optimization problem (2) is always realized (Theorem 26).
  • Several geometric properties of the metric space $(\mathcal{M}_{\sim}^{Z,p},\mathrm{GW}_{p}^{Z})$ are established: it is separable (Proposition 36), complete if and only if $Z$ is (Theorem 39), path-connected (Theorem 40), and geodesic if $Z$ is (Theorem 41). Many of these results are novel when restricted to the examples of $Z$-GW distances covered in Theorem 12.
@@
The structure of the paper is as follows. After concluding the introduction section with a discussion of related work, the main definitions of the objects of interest are given in Section 2. Section 3 describes examples of $Z$-GW distances, including some distances which have already appeared in the literature and some which are novel to the present paper. The basic geometric properties of $Z$-GW distances are established in Section 4. Finally, the paper concludes with our results on approximations of $Z$-GW distances in Section 5.
@@ 2.1.1 Optimal Transport Distances
Definition 2 (Wasserstein Distance).
@@
Definition 3 (Measure Network [19]).
@@
Definition 5 (Gromov-Wasserstein Distance [44, 19]).
@@
In this article we will introduce a generalization of the GW distance and when clarity is needed, we refer to $\mathrm{GW}_{p}$ as the standard Gromov-Wasserstein (GW) $p$-distance. The (standard) GW $p$-distance is finite when restricted to the space of measure networks whose network kernels have finite $p$-th moment. On this subspace, it defines a pseudometric whose distance-zero equivalence classes can be precisely characterized: see [19, Theorem 2.4] or the discussion in Section 4 below.
@@ 2.1.2 Metric Space-Valued $L^{p}$-Spaces
The main object of study in this paper generalizes the above definitions to the case where the network kernel $\omega$ takes values in a general metric space. To ensure that the distance is finite, we will assume that the network kernel is an $L^{p}$ function, in the sense of [38, Section 1.1]. This parallels the definition of the Wasserstein distance, where we assume that the measures have finite $p$-th moments.
We will now recall the definition of $L^{p}$ functions valued in a metric space. For measurable functions $f$ and $g$, the natural definition of a distance between $f$ and $g$ is the integral of $d(f(x),g(x))$, but this function is not always measurable due to Nedoma’s pathology [61, Section 15.9]. However, if we assume that the target metric space is separable, the distance is measurable, and the integral is well-defined. With this in mind, we present the following definition.
Definition 6 (Metric Space-Valued $L^{p}$ Spaces [38, Section 1.1]).
@@
Remark 9.
Item 2 and Item 3 in Proposition 8 are classical results when $Y$ is a Banach space. However, to the best of our knowledge, the case when $Y$ is a general separable metric space has not been studied in previous literature.
Proof of Proposition 8.
@@
    The main idea of the proof is due to [29]. By the theorem of Banach [5, Chapter XI, §$\S$8, Theorem 10], $Y$ can be isometrically embedded onto a subset of a separable Banach space $C([0,1])$ of real-valued continuous functions on $[0,1]$ with the sup-norm. By Pettis measurability theorem [30, Theorem 1.1.6], all measurable functions with values in $Y$ are strongly measurable by separability of $Y$, so $L^{p}(X,\mu_{X};Y)$ is a subset (or a closed subspace) of the Lebesgue-Bochner space $L^{p}(X,\mu_{X};C([0,1]))$. Since $X$ is separable, its Borel $\sigma$-algebra is countably generated, so $L^{p}(X,\mu_{X};C([0,1]))$ is separable [30, Theorem 1.2.29]. Therefore, $L^{p}(X,\mu_{X};Y)$ is separable.
@@ 2.2 $Z$-Valued Measure Networks
In this subsection, we introduce our main objects of study, which generalize the notion of a measure network (Definition 3) and an extension of GW distance for comparing them. The basic idea is to allow network kernels to take values in some fixed, but arbitrary metric space. Similar ideas were considered recently in the machine learning literature [78, 34], with applications to analysis of attributed graph datasets, but the idea goes back at least to [31]—see the discussion in Section 1.2. The definitions given below generalize those which have appeared in the previous literature.
@@ 2.2.1 Main Definitions
Definition 11 (Gromov-Wasserstein Distance Between $(Z,p)$-Networks).
@@ 3 Examples of $Z$-Networks and $Z$-Gromov-Wasserstein Distances
We show below in Section 4 that the $Z$-Gromov-Wasserstein distance of Definition 11 induces a metric on a certain quotient of the space of $Z$-networks, and establish some of its geometric properties. To motivate this, we first provide many examples of $Z$-GW distances; these include metrics which have already appeared in the literature (Section 3.1) and other natural metrics that appear to be novel (Section 3.2). An important takeaway message is that it is not always necessary to re-establish metric properties of variants of GW distances from scratch—a practice which is common in the recent literature—since these properties frequently follow immediately from high-level principles.
@@ 3.1 Examples of existing $Z$-Gromov-Wasserstein Distances in the Literature
The definitions of these optimal transport-type distances are provided below, as necessary, and the proof of the theorem is split among Propositions 13, 14, 15, 16, 23, 24, and 25.
@@ 3.1.1 Wasserstein distance and Gromov-Wasserstein distances
Proposition 13.
The standard Gromov-Wasserstein distance between measure networks (Definition 5) is a $Z$-GW distance with $(Z,d_{Z})$ equal to $\mathbb{R}$ with its standard metric.
@@
but equality does not hold in general—see [53, Theorem 5.1(c)]. Next, we show that $Z$-GW distance does give such a generalization.
Proposition 14.
The Wasserstein distance over an arbitrary metric space $(Z,d_{Z})$ (Definition 2) can be realized as a $Z$-GW distance.
@@ 3.1.2 Attributed Graphs
Definition 18 (Attributed Network).
@@
Remark 19.
The definition above is more general than the one given in [78, Definition 2.1]: therein, all functions are assumed to be bounded and continuous. Moreover, [78] does not assume separability of the target metric spaces, which can potentially lead to technical issues of well-definededness, following the discussion in Section 2.1.2.
@@
Definition 20 (Fused Network Gromov-Wasserstein Distnace).
@@
Remark 21.
The formulation given in (16) is actually subtly different than the one given in [78]. In analogy with Remark 17, we have added an extra power of $1/q$ to the integrand. We will see below that this leads to improved theoretical properties of the distance.
@@
    Taking $\alpha=1$, $\beta=0$, we recover the $Z$-GW $p$-distance (Definition 11) between the underlying $Z$-networks, with $(Z,d_{Z})=(\Omega,d_{\Omega})$.
@@
Remark 22.
The original formulation of FGW distance also omitted the $q$-th root in the integrand that was discussed in Remark 21, but we argue for its inclusion for theoretical reasons discussed below. The FGW framework introduced in [72] is also slightly more general, in that it considers a joint distribution on $X\times\Psi$, rather than a function $\psi_{X}:X\to\Psi$; the formulation considered here is the one most commonly used in practice—see [70, 23].
@@ 3.1.3 Diffusions and Stochastic Processes
Inspired on a notion of spectral convergence for Riemannian manifods introuced by Kasue and Kumura [33], in [45, 46] the author considers a certain spectral version of the GW distance between Riemannian manifolds (see S. Lim’s PhD thesis, [42, Chapter 5], for a generalization to Markov processes). For a given compact Riemannian manifold $(M,g_{M})$ let $k_{M}:M\times M\times\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ denote its normalized heat kernel and $\mu_{M}$ its normalized volume measure (we use $\mathbb{R}_{>0}$ to denote the set of positive real numbers). Then, if $N$ is another compact Riemannian manifold and $p\geq 1$, the spectral Gromov-Wasserstein $p$-distance between $M$ and $N$ is
@@
To recast $\mathrm{GW}_{p}^{\mathrm{spec}}$ as a $Z$-GW distance, we make the following choices:
@@
We represent a compact Riemannian manifold $(M,g_{M})$ as a $Z$-measure network $(M,\omega_{M},\mu_{M})$ with $\mu_{M}$ as above and $\omega_{M}:M\times M\to\mathbb{R}$ defined by
@@
$\inf_{\pi\in\mathcal{C}(\mu_{X},\mu_{Y})}\|d_{Z}\circ(\omega_{M}\otimes\omega_{N})\|_{L^{p}(\pi\times\pi)}$ $=\inf_{\pi\in\mathcal{C}(\mu_{X},\mu_{Y})}\left\|\sup_{t>0}c^{2}(t)\Gamma^{\mathrm{spec}}_{M,N,t}\right\|_{L^{p}(\pi\otimes\pi)}$ (24)
$=\inf_{\pi\in\mathcal{C}(\mu_{X},\mu_{Y})}\sup_{t>0}c^{2}(t)\|\Gamma^{\mathrm{spec}}_{M,N,t}\|_{L^{p}(\pi\otimes\pi)}.$ (25)
This proves the following result.
Proposition 24.
The spectral Gromov-Wasserstein distance of [45] is a $Z$-GW distance with $(Z,d_{Z})$ as in (20) and (21).

### v2

@@ Abstract
The Gromov-Wasserstein (GW) distance is a powerful tool for comparing metric measure spaces which has found broad applications in data science and machine learning. Driven by the need to analyze datasets whose objects have increasingly complex structure (such as node and edge-attributed graphs), several variants of GW distance have been introduced in the recent literature. With a view toward establishing a general framework for the theory of GW-like distances, this paper considers a vast generalization of the notion of a metric measure space: for an arbitrary metric space $Z$, we define a $Z$-network to be a measure space endowed with a kernel valued in $Z$. We introduce a method for comparing $Z$-networks by defining a generalization of GW distance, which we refer to as $Z$-Gromov-Wasserstein ($Z$-GW) distance. This construction subsumes many previously known metrics and offers a unified approach to understanding their shared properties. This paper demonstrates that the $Z$-GW distance defines a metric on the space of $Z$-networks which retains desirable properties of $Z$, such as separability, completeness, and geodesicity. Many of these properties were unknown for existing variants of GW distance that fall under our framework. Our focus is on foundational theory, but our results also include computable lower bounds and approximations of the distance which will be useful for practical applications.
@@ 1 Introduction
where the infimum is over probability measures $\pi$ on $X\times Y$ whose left and right marginals are $\mu_{X}$ and $\mu_{Y}$, respectively. Intuitively, such a measure describes a probabilistic correspondence between the sets $X$ and $Y$, and the integral in (1) measures the extent to which such a correspondence distorts the metric structures of the spaces. The optimization problem (1) therefore seeks a probabilistic correspondence which minimizes total metric distortion. It was shown in [52] that the Gromov-Wasserstein $p$-distance defines a metric on the space of all triples $(X,d_{X},\mu_{X})$, considered up to measure-preserving isometries, when $(X,d_{X})$ is compact and $\mu_{X}$ is fully supported (the compactness assumption can be relaxed [57, 77]).
The Gromov-Wasserstein (GW) framework has become a popular tool in data science and machine learning—see [29, 25, 16, 88, 24, 1], among many others—due to its flexibility in handling diverse data types, its robustness to changes in metric or measure structures [62, Theorem 5.1], and recent advances in scalable and empirically accurate computational schemes for its estimation [65, 89, 22, 70, 48, 85]. It was observed in [65] that the formula for GW distance in (1) still makes sense when the condition that $d_{X}$ and $d_{Y}$ are metrics is relaxed; that is, the formula gives a meaningful comparison between structures of the form $(X,\omega_{X},\mu_{X})$, where $\omega_{X}:X\times X\to\mathbb{R}$ is an arbitrary measurable function. This point of view was later studied formally in [23], where it was shown that the GW distance defines a metric on the space of such triples $(X,\omega_{X},\mu_{X})$, considered up to a natural notion of equivalence. This is convenient, for example, when handling graph datasets, where it is natural to represent a graph’s structure through its adjacency function, Laplacian, or heat kernel.
Driven by applications to machine learning on increasingly complex data types, several variants of GW distance have been introduced in the literature [53, 84, 43, 56, 2, 90]. For example, [84] introduces an adaptation of GW distance which is equipped to handle graphs endowed with node features; that is, graphs whose nodes are endowed with values in some auxiliary metric space (in fact, this idea goes back further to, at least, [17, Section 5], where a similar construction provided variants of both the Gromov-Hausdorff distance and the $p=\infty$ verison of GW distance). Each time that a new variant of GW distance is introduced in the literature, its metric properties have been re-established. Overviewing these proofs reveals that they tend to follow a common template, suggesting the existence of a higher-level explanation of their shared properties.
@@ 1.1 Main Contributions and Outline
  • Theorem 3 says that, when $(Z,d_{Z})$ is separable, the $Z$-GW distance $\mathrm{GW}_{p}^{Z}$ defines a metric on the space $\mathcal{M}_{\sim}^{Z,p}$ consisting of $Z$-networks, considered up to a natural notion of equivalence (Definition 4.2). As a consequence, Corollary 4.3 shows that Fused GW and Fused Network GW distances are metrics; these were previously only shown to satisfy a certain weak triangle inequality. The proof of Theorem 3 relies on a technical result, which says that the solution of the optimization problem (2) is always realized (Theorem 2).
  • Several geometric and topological properties of the metric space $(\mathcal{M}_{\sim}^{Z,p},\mathrm{GW}_{p}^{Z})$ are established: it is separable (Proposition 4.8), complete if and only if $Z$ is (Theorem 5), contractible, regardless of the topology of $Z$ (Theorem 6), and geodesic if $Z$ is (Theorem 7). Many of these results are novel when restricted to the examples of $Z$-GW distances covered in Theorem 1.
@@
The structure of the paper is as follows. After concluding the introduction section with a discussion of related work, the main definitions of the objects of interest are given in Section 2. Section 3 describes examples of $Z$-GW distances, including some distances which have already appeared in the literature and some which are novel to the present paper. The basic geometric properties of $Z$-GW distances are established in Section 4. Finally, our results on approximations of $Z$-GW distances are presented in Section 5. The paper concludes with a discussion of open questions and future research directions in Section 6.
@@ 1.2 Related Work
Another recent approach to a general framework for Gromov-Wasserstein-like distances is developed in [91]. In that paper, the generalization is given by allowing optimization over spaces of measure couplings with a certain additional structure, rather than varying the target of the kernel function. That framework encompasses a different collection of variants of GW distances (namely, those of the form presented in [26, 81]) than those considered here. Moreover, [91] is mainly focused on curvature bounds and on a computational framework, rather than the structural results presented in this paper.
@@ 2.1.1 Optimal Transport Distances
Definition 2.2 (Wasserstein Distance).
@@
Definition 2.3 (Measure Network [23]).
@@
Definition 2.5 (Gromov-Wasserstein Distance [52, 23]).
@@
In this article we will introduce a generalization of the GW distance and when clarity is needed, we refer to $\mathrm{GW}_{p}$ as the standard Gromov-Wasserstein (GW) $p$-distance. The (standard) GW $p$-distance is finite when restricted to the space of measure networks whose network kernels have finite $p$-th moment. On this subspace, it defines a pseudometric whose distance-zero equivalence classes can be precisely characterized: see [23, Theorem 2.4] or the discussion in Section 4 below.
@@ 2.1.2 Metric Space-Valued $L^{p}$-Spaces
The main object of study in this paper generalizes the above definitions to the case where the network kernel $\omega$ takes values in a general metric space. To ensure that the distance is finite, we will assume that the network kernel is an $L^{p}$ function, in the sense of [46, Section 1.1]. This parallels the definition of the Wasserstein distance, where we assume that the measures have finite $p$-th moments.
We will now recall the definition of $L^{p}$ functions valued in a metric space. For measurable functions $f$ and $g$, the natural definition of a distance between $f$ and $g$ is the integral of $d(f(x),g(x))$, but this function is not always measurable due to Nedoma’s pathology [71, Section 15.9]. However, if we assume that the target metric space is separable, the distance is measurable, and the integral is well-defined. With this in mind, we present the following definition.
Definition 2.6 (Metric Space-Valued $L^{p}$ Spaces [46, Section 1.1]).
@@
Remark 2.9.
Item 2 and Item 3 in Proposition 2.8 are classical results when $Y$ is a Banach space. However, to the best of our knowledge, the case when $Y$ is a general separable metric space has not been studied in previous literature.
Proof of Proposition 2.8.
@@
    The main idea of the proof is due to [37]. By the theorem of Banach [7, Chapter XI, §$\S$8, Theorem 10], $Y$ can be isometrically embedded onto a subset of a separable Banach space $C([0,1])$ of real-valued continuous functions on $[0,1]$ with the sup-norm. By Pettis measurability theorem [38, Theorem 1.1.6], all measurable functions with values in $Y$ are strongly measurable by separability of $Y$, so $L^{p}(X,\mu_{X};Y)$ is a subset (or a closed subspace) of the Lebesgue-Bochner space $L^{p}(X,\mu_{X};C([0,1]))$. Since $X$ is separable, its Borel $\sigma$-algebra is countably generated, so $L^{p}(X,\mu_{X};C([0,1]))$ is separable [38, Theorem 1.2.29]. Therefore, $L^{p}(X,\mu_{X};Y)$ is separable.
@@ 2.2 $Z$-Valued Measure Networks
In this subsection, we introduce our main objects of study, which generalize the notion of a measure network (Definition 2.3) and an extension of GW distance for comparing them. The basic idea is to allow network kernels to take values in some fixed, but arbitrary metric space. Similar ideas were considered recently in the machine learning literature [90, 42], with applications to analysis of attributed graph datasets, but the idea goes back at least to [39]—see the discussion in Section 1.2. The definitions given below generalize those which have appeared in the previous literature.
@@ 2.2.1 Main Definitions
Definition 2.11 (Gromov-Wasserstein Distance Between $(Z,p)$-Networks).
@@ 3 Examples of $Z$-Networks and $Z$-Gromov-Wasserstein Distances
We show below in Section 4 that the $Z$-Gromov-Wasserstein distance of Definition 2.11 induces a metric on a certain quotient of the space of $Z$-networks, and establish some of its geometric properties. To motivate this, we first provide many examples of $Z$-GW distances; these include metrics which have already appeared in the literature (Section 3.1) and other natural metrics that appear to be novel (Section 3.2). An important takeaway message is that it is not always necessary to re-establish metric properties of variants of GW distances from scratch—a practice which is common in the recent literature—since these properties frequently follow immediately from high-level principles.
@@ 3.1 Examples of existing $Z$-Gromov-Wasserstein Distances in the Literature
The definitions of these optimal transport-type distances are provided below, as necessary, and the proof of the theorem is split among Propositions 3.1, 3.2, 3.3, 3.4, 3.11, 3.12, and 3.13.
@@ 3.1.1 Wasserstein distance and Gromov-Wasserstein distances
Proposition 3.1.
The standard Gromov-Wasserstein distance between measure networks (Definition 2.5) is a $Z$-GW distance with $(Z,d_{Z})$ equal to $\mathbb{R}$ with its standard metric.
@@
but equality does not hold in general—see [62, Theorem 5.1(c)]. Next, we show that $Z$-GW distance does give such a generalization.
Proposition 3.2.
The Wasserstein distance over an arbitrary metric space $(Z,d_{Z})$ (Definition 2.2) can be realized as a $Z$-GW distance.
@@ 3.1.2 Attributed Graphs
Definition 3.6 (Attributed Network).
@@
Remark 3.7.
The definition above is more general than the one given in [90, Definition 2.1]: therein, all functions are assumed to be bounded and continuous. Moreover, [90] does not assume separability of the target metric spaces, which can potentially lead to technical issues of well-definededness, following the discussion in Section 2.1.2.
@@
Definition 3.8 (Fused Network Gromov-Wasserstein Distnace).
@@
Remark 3.9.
The formulation given in (16) is actually subtly different than the one given in [90]. In analogy with Remark 3.5, we have added an extra power of $1/q$ to the integrand. We will see below that this leads to improved theoretical properties of the distance.
@@
    Taking $\alpha=1$, $\beta=0$, we recover the $Z$-GW $p$-distance (Definition 2.11) between the underlying $Z$-networks, with $(Z,d_{Z})=(\Omega,d_{\Omega})$.
@@
Remark 3.10.
The original formulation of FGW distance also omitted the $q$-th root in the integrand that was discussed in Remark 3.9, but we argue for its inclusion for theoretical reasons discussed below. The FGW framework introduced in [84] is also slightly more general, in that it considers a joint distribution on $X\times\Psi$, rather than a function $\psi_{X}:X\to\Psi$; the formulation considered here is the one most commonly used in practice—see [80, 32].
@@ 3.1.3 Diffusions and Stochastic Processes
Inspired on a notion of spectral convergence for Riemannian manifods introuced by Kasue and Kumura [41], in [53, 54] the author considers a certain spectral version of the GW distance between Riemannian manifolds (see S. Lim’s PhD thesis, [50, Chapter 5], for a generalization to Markov processes). For a given compact Riemannian manifold $(M,g_{M})$ let $k_{M}:M\times M\times\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ denote its normalized heat kernel and $\mu_{M}$ its normalized volume measure (we use $\mathbb{R}_{>0}$ to denote the set of positive real numbers). Then, if $N$ is another compact Riemannian manifold and $p\geq 1$, the spectral Gromov-Wasserstein $p$-distance between $M$ and $N$ is
@@
To recast $\mathrm{GW}_{p}^{\mathrm{spec}}$ (or, rather, a variant thereof—see below) as a $Z$-GW distance, we make the following choices:
@@
We represent a compact Riemannian manifold $(M,g_{M})$ as a $Z$-measure network $(M,\omega_{M},\mu_{M})$ with $\mu_{M}$ as above and $\omega_{M}:M\times M\to Z$ defined by
@@
$\|d_{Z}\circ(\omega_{M}\times\omega_{N})\|_{L^{p}(\pi\times\pi)}$ $=\left\|\sup_{t>0}c^{2}(t)\Gamma^{\mathrm{spec}}_{M,N,t}\right\|_{L^{p}(\pi\otimes\pi)}.$ (24)
Accordingly, we define a variant of the spectral GW $p$-distance:
$\widetilde{\mathrm{GW}}_{p}^{\mathrm{spec}}(M,N):=\frac{1}{2}\inf_{\pi\in\mathcal{C}(\mu_{X},\mu_{Y})}\left\|\sup_{t>0}c^{2}(t)\Gamma^{\mathrm{spec}}_{M,N,t}\right\|_{L^{p}(\pi\otimes\pi)}.$
We have the following result.
Proposition 3.12.
The spectral GW $p$-distance $\widetilde{\mathrm{GW}}_{p}^{\mathrm{spec}}$ is a $Z$-GW distance, with $(Z,d_{Z})$ as in (20) and (21), which upper bounds the spectral Gromov-Wasserstein distance of [53].
Proof.
The fact that $\widetilde{\mathrm{GW}}_{p}^{\mathrm{spec}}$ is a $Z$-GW distance follows immediately by definition. To verify the upper bound claim, let $M$ and $N$ be compact Riemannian manifolds and choose an arbitrary coupling $\pi$ between their normalized volume measures. We then consider the map
$F:(X\times Y\times X\times Y)\times\mathbb{R}_{>0}$ $\to\mathbb{R}$ (25)
$\big{(}(x,y,x^{\prime},y^{\prime}),t\big{)}$ $\mapsto c^{2}(t)\Gamma^{\mathrm{spec}}_{M,N,t}(x,y,x^{\prime},y^{\prime})$ (26)
as a function on the product of measure spaces $(X\times Y\times X\times Y,\pi\otimes\pi)$ and $(\mathbb{R},\mathcal{L})$, where we use $\mathcal{L}$ to denote Lebesgue measure on $\mathbb{R}_{>0}$ through the duration of the proof. We now compare the objective functions of the optimization problems associated to the distances $\widetilde{\mathrm{GW}}_{p}^{\mathrm{spec}}(M,N)$ and $\mathrm{GW}_{p}^{\mathrm{spec}}(M,N)$ as
$\left\|\sup_{t>0}c^{2}(t)\Gamma^{\mathrm{spec}}_{M,N,t}\right\|_{L^{p}(\pi\otimes\pi)}$ $=\left\|\left\|F\right\|_{L^{\infty}(\mathcal{L})}\right\|_{L^{p}(\pi\otimes\pi)}$ (27)
$\geq\left\|\left\|F\right\|_{L^{p}(\pi\otimes\pi)}\right\|_{L^{\infty}(\mathcal{L})}$ (28)
$=\sup_{t>0}\|c^{2}(t)\Gamma_{M,N,t}^{\mathrm{spec}}\|_{L^{p}(\pi\otimes\pi)}$ (29)
$=\sup_{t>0}c^{2}(t)\cdot\|\Gamma_{M,N,t}^{\mathrm{spec}}\|_{L^{p}(\pi\otimes\pi)},$ (30)
where we have interchanged suprema and $L^{\infty}$ norms via continuity in $t$, and we have applied a generalized version of Minkowski’s inequality [4, Proposition 1.3]. Since the estimate holds for an arbitrary coupling $\pi$, this shows that $\widetilde{\mathrm{GW}}_{p}^{\mathrm{spec}}(M,N)\geq\mathrm{GW}_{p}^{\mathrm{spec}}(M,N)$. ∎